1. With the following information Setup the balance sheet for Eagle Bank when the reserve requirement is 10%: Equity 20, Securities 70, Savings deposits 10, Personal Loans 40, Retained earnings 20, Time deposits 30, Home mortgages 60, Demand deposits 100.

2. Show adjustments in the Eagle balance sheet when there is a loan default of 10.

3. Show adjustments in the Eagle balance sheet when there is a deposit withdrawal of 40.

4. Show adjustments in the Eagle balance sheet when equity is increased by 10.
Answer #1

1. Solution:

Assets will be those entries which the bank will have to receive back from someone.

Liabilities are those which are going to be paid back to someone by the bank.

reserve requirements are those cash which has to be kept with the central bank from the net time, demand and savings deposit. hence our total deposit here is 140 hence 14 of those will be with the central bank.

hence cash left with bank= (140-14)+retaining earnings+Equity+Home mortgage-Personal loans-securities.

and according to balance sheet principle, Assets are always equal to the liabilities.

Assets

Liabilities

Personal Loans

40

Demand deposits

100

Securities

70

Savings deposits

10

Cash reserves with the central bank(10% Reserve Ratio)

14

Time deposits

30

Cash with the bank

116

Retained earnings

20

Home mortgages

60

Equity

20

total 240 total 240

2. Solution:

When loan default occurs we have to write off the loan and add it as a non-performing asset. hence personal loan asset is reduced by 10 and NPA increased by 10.

Assets

Liabilities

Personal Loans

40-10=30

Demand deposits

100

Securities

70

Savings deposits

10

Cash reserves with central bank(10% Reserve Ratio)

14

Time deposits

30

Cash with the bank

116

Retained earnings

20

Loan default(Non performing Asset)

10

Home mortgages

60

Equity

20

total 240 total 240

3. solution:

when demand deposit gets a withdrawal of 40. then demand deposits get reduced to 60.

but now total reserve requirements also get reduced by 10% of 40 which is 4.

now a total reserve with the central bank becomes 14-4=10

and total cash with the bank increases by 4 and reduced by 40 and becomes 80.

Assets

Liabilities

Personal Loans

40

Demand deposits

100-40=60

Securities

70

Savings deposits

10

Cash reserves with central bank(10% Reserve Ratio)

14-4=10

Time deposits

30

Cash with the bank

116-40+4=80

Retained earnings

20

Home mortgages

60

Equity

20

total 200 total 200

4. Solution:

equity gets increased by 10 so liability increases by 10. now when selling the equity shares banks also gets 10 extra cash amount. hence cash becomes 126.

Assets

Liabilities

Personal Loans

40

Demand deposits

100

Securities

70

Savings deposits

10

Cash reserves with central bank(10% Reserve Ratio)

14

Time deposits

30

Cash with the bank

116+10=126

Retained earnings

20

Home mortgages

60

Equity

20+10

total

250

total

250

I hope you liked the answer. Have a great day ahead. Comment if you have any doubt. I'll get back to you ASAP.

thanks.

give good feedback :)

study hard.